OpenStudy (anonymous):

Solve for x. c-1/x + d= cd

OpenStudy (kc_kennylau):

\(\dfrac{c-1}x+d=cd\)?

OpenStudy (anonymous):

yes

OpenStudy (australopithecus):

multiply both sides by x first off

OpenStudy (shamil98):

\[\huge \frac{ c-1 }{ x } = cd - d\] \[\huge \frac{ c-1 }{ cd-d}= x\]

OpenStudy (kc_kennylau):

c-1+xd=cxd c-1=cxd-xd c-1=x(cd-d) x=(c-1)/(cd-d)

OpenStudy (kc_kennylau):

\[\begin{array}{rl}\frac{c-1}x+d&=&cd\\c-1+xd&=&cxd\\cxd-xd&=&c-1\\x(cd-d)&=&c-1\\x&=&\frac{c-1}{xd-d}\end{array}\]

OpenStudy (australopithecus):

kc_kennylau you have posted the best answer but you shouldnt leave steps out

OpenStudy (australopithecus):

it makes it harder for the original poster to follow your solution, never make assumptions

OpenStudy (anonymous):

would it be C over D?

OpenStudy (kc_kennylau):

sorry i meant \(x=\dfrac{c-1}{cd-d}\)

OpenStudy (australopithecus):

\[\frac{c-1}{x} + d = cd\] \[x(\frac{c-1}{x} + d) = x(cd)\] \[\frac{x}{1}(\frac{c-1}{x} + d) = x(cd)\] \[\frac{x*d}{1} + \frac{(c-1)x}{1*x} = xcd\] \[x*d + \frac{(c-1)x}{x} = xcd\] \[x*d + (c-1)*1 = xcd\] \[x*d + (c-1) = xcd\] \[(c-1) = xcd - xd\] \[(c-1) = (cd - d)x\] \[\frac{(c-1)}{(cd - d)} = \frac{(cd - d)x}{(cd - d)}\] \[\frac{(c-1)}{(cd - d)} = 1*x\] \[\frac{(c-1)}{(cd - d)} =x\]

OpenStudy (australopithecus):

remember, \[\frac{x}{1} = x\]

OpenStudy (australopithecus):

you can divide anything by 1 and it wont change its value

OpenStudy (kc_kennylau):

wow @Australopithecus you have done a great job

OpenStudy (kc_kennylau):

did you type out all the \(\LaTeX\) yourself or you used the equation button

OpenStudy (australopithecus):

well I did leave some steps out

OpenStudy (australopithecus):

I used the equation button

OpenStudy (anonymous):

Thanks you guys helped me out a lot.

OpenStudy (australopithecus):

like I should have written \[\frac{d}{1}\] instead of d

OpenStudy (australopithecus):

but I think he got it at least I hope these are fundamental rules